OpenStudy (jenessa_2413):

A box contains 100 colored chips; some are yellow and some are green. John chooses a chip at random, records the color, and places it back in the bag. John has recorded 57 yellow chips and 19 green chips. Using these results, what is the predicted number of green chips in the box? 25 43 76 81

OpenStudy (chrishogan):

Give me a minute, I'll help you. :)

OpenStudy (jenessa_2413):

ok thanks

OpenStudy (chrishogan):

Have you done problems like this in the past?

OpenStudy (jenessa_2413):

yea but not a lot

OpenStudy (jenessa_2413):

u there???

OpenStudy (chrishogan):

Yep just doing the math.

OpenStudy (chrishogan):

lets say we actually have 10 yellow and 30 green our ratio is 1 to 3 1k + 3k = 40, what is k?

OpenStudy (jenessa_2413):

0?

OpenStudy (chrishogan):

Not quite, we need 1(x) + 3(x) to equal 40

OpenStudy (chrishogan):

I'll just give you this one. add your common variables (1x,3x) \[1x+3x=40\] Divide both sides by 4 to find your x \[4x=40\] \[\frac{ 4x }{ 4 }=\frac{ 40 }{ 4 }\]

OpenStudy (chrishogan):

so what is x? :)